40m
QIL
Cryo_Lab
CTN
SUS_Lab
CAML
OMC_Lab
CRIME_Lab
FEA
ENG_Labs
OptContFac
Mariner
WBEEShop
|
40m Log |
Not logged in |
 |
|
Wed Jan 25 15:41:29 2017, Lydia, Update, IMC, 29.5 MHz modulation depth measurement plan
|
Wed Jan 25 17:30:03 2017, Koji, Update, IMC, 29.5 MHz modulation depth measurement plan
|
Wed Jan 25 19:39:07 2017, gautam , Update, IMC, 29.5 MHz modulation depth measurement plan
|
Fri Jan 27 21:25:11 2017, Lydia, Update, IMC, 29.5 MHz modulation depth
|
Sat Jan 28 01:25:51 2017, gautam, Update, IMC, 29.5 MHz modulation depth 
|
Mon Jan 30 19:07:48 2017, gautam, Update, IMC, RF AM stabilization box pulled out
|
Tue Jan 31 01:07:20 2017, Lydia, Update, IMC, RF AM stabilization box pulled out
|
Tue Jan 31 13:46:34 2017, rana, Update, IMC, RF AM stabilization box pulled out
|
Tue Jan 31 14:17:48 2017, gautam, Update, IMC, RF AM stabilization box pulled out
|
Tue Jan 31 22:07:13 2017, gautam, Update, IMC, RF AM stabilization box revamp
|
Tue Jan 31 22:28:39 2017, Lydia, Update, IMC, RF AM stabilization box pulled out
|
Wed Feb 1 11:51:19 2017, Koji, Update, IMC, RF AM stabilization box pulled out
|
Wed Feb 1 16:45:56 2017, Lydia, Update, IMC, RF AM stabilizer box Modification Plan
|
Wed Feb 1 20:49:34 2017, rana, Update, IMC, RF AM stabilizer box Modification Plan
|
Wed Feb 1 23:13:30 2017, Lydia, Update, IMC, RF AM stabilizer box Modification Plan
|
Sun Feb 5 21:56:50 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement
|
Tue Feb 7 12:01:10 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement
|
Tue Feb 7 17:00:55 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement
|
Fri Feb 10 02:14:10 2017, gautam, Update, IMC, 29.5 MHz stabilizer box replacement
|
Fri Feb 10 11:41:43 2017, Lydia, Update, IMC, 29.5 MHz stabilizer box replacement
|
Fri Feb 10 13:24:28 2017, rana, Update, IMC, 29.5 MHz stabilizer box replacement
|
Fri Feb 10 18:21:21 2017, gautam, Update, IMC, IMC Demod board
|
Fri Feb 10 19:32:15 2017, Koji, Update, IMC, IMC Demod board
|
Sun Feb 12 01:16:57 2017, gautam, Update, IMC, IMC length loop - summary of changes    
|
Mon Feb 13 11:55:14 2017, rana, Update, IMC, IMC length loop - summary of changes
|
Mon Feb 13 13:34:44 2017, gautam, Update, IMC, IMC length loop - bad SMA cable replaced
|
Mon Feb 13 19:44:55 2017, Lydia, Update, IMC, Front panel for 29.5 MHz amplifier box
|
Wed Mar 1 21:15:40 2017, Lydia, Update, IMC, Front panel for 29.5 MHz amplifier box
|
Wed Mar 1 23:56:09 2017, gautam, Update, IMC, Front panel for 29.5 MHz amplifier box
|
Thu Mar 2 20:32:18 2017, Lydia, Update, IMC, Front panel for 29.5 MHz amplifier box
|
Wed Feb 15 23:54:13 2017, gautam, Update, IMC, IMC saga continues... 
|
Fri Feb 17 20:10:18 2017, gautam, Update, IMC, WFS servos turned back on
|
Sat Feb 18 14:09:06 2017, rana, Update, IMC, WFS servos turned back on
|
Sat Feb 18 21:50:48 2017, gautam, Update, IMC, WFS servos turned back on
|
Sun Mar 5 12:41:23 2017, gautam, Update, IMC, WFS servo-steppin  
|
|
Message ID: 12822
Entry time: Sun Feb 12 01:16:57 2017
In reply to: 12821
Reply to this: 12823
12824
|
Author: |
gautam |
Type: |
Update |
Category: |
IMC |
Subject: |
IMC length loop - summary of changes |
|
|
29.5 MHz RF Modulation Source
- The +13dBm from the Wenzel oscillator gets amplified to +27dBm by a ZHL-2-S. There is a 5dB attenuator on the input to the amplifier to avoid compression/saturation.
- The amplified output goes to the EOM (+26dBm measured at the rack, no measurement done at the input to the triple-resonant circuit box yet), while a 10dB coupled part goes to the RF distribution box which splits the input into 16 equal parts. The outputs were measured to spit out +5dBm.
- 2 of these go to the WFS demod boards - it was verified that this level of drive is okay for the comparator chips on the demod board.
- A third output goes to the IMC Demod board. The demod board was modified so that the nominal LO input level is now +5dBm (details below).
- The remaining outputs are all terminated with 50ohms.
IMC Demodulation Board
- The input attenuator, amplifier and power splitter were removed.
- Schematic with changes marked and power levels measured, along with a high-res photograph (taken with our fancy new Macro lens + LED light ring) has been uploaded to a page I made to track changes for this part on the DCC (linked to 40m document tree).
- After making the changes, it was verified that the power levels in the signal chain were appropriate up till the input to the ERA-5SM amplifier directly before the LO. These levels were deemed appropriate, and also scaled in a predictable manner with the input power. As Koji mentions in the previous elog, the dynamically changing input impedance of the mixer makes it difficult to measure the LO level at this point, but I am satisfied that it is within ~1dBm of the nominal +17dBm the mixer wants.
- The board was further checked for gain imbalance and orthogonality of the I and Q outputs. The graphic below show that there is negligible gain imbalance, but the relative phase between the I and Q channels is ~78 degrees (they should be 90 degrees). Of course this doesn't matter for the IMC locking as we only use the I phase signal, but presumably, we want to understand this effect and compensate for it.

- The label on the front panel has been updated to reflect the fact that the nominal LO input is now +5dBm
- The demodulation phase had changed since the RF signal change was modified - Rana and I investigated this effect on Monday morning, and found that a new ~1.5m long cable was needed to route the signal from the RF distribution box to the LO input of the demod board, which I made. Subsequent modifications on the demod board meant that an extra ~10cm length was needed, so I just tacked on a short length of cable. All of the demodulated signal is now in the I output of the demod board (whereas we had been using the Q output).
- The graphics below confirms that claim above. Note the cool feature on the digital scopes that the display persistence can be set to "infinity"!

I wanted to do a quick check to see if the observed signal levels were in agreement with tests done on the workbench with the Marconi. The mixers used, JMS-1H, have an advertised conversion loss of ~7dB (may be a little higher if we are not driving the LO at +17dBm). The Lissajous ellipse above is consistent with these values. I didn't measure powers with the MC REFL PD plugged into the demod board, but the time series plot above suggest that I should have ~0dBm power in the MC REFL PD signal at 29.5MHz for the strongest flashes (~0.3Vpp IF signal for the strong flashes).
MC Servo Board
- As mentioned above, we now use the I phase signal for lMC PDH locking.
- This has resulted in an overall sign change of the servo. I have updated the MEDM screen to reflect that "MINUS" is the correct polarity now..
- To set the various gains, I measured the OLTF for various configurations using the usual IN1/IN2 prescription on the MC Servo Board (using the Agilent analyzer).
- I started at 0dBm "In1 Gain", and the nominal (old) values for "VCO gain", "FSS Common Gain" and "FSS FAST gain" and found that though I could lock the MC, I couldn't reliably turn on the boosts.
- After some tweaking, I settled on +10dB "In1 Gain". Here, locking was much more reliable, and I was able to smoothly turn on the Super Boosts. The attached OLTF measurement suggests a UGF of ~118kHz and phase margin of a little more than 30 degrees. There is room for optimization here, since we have had UGFs closer to 200kHz in the recent past.

- I didn't get around to measuring the actual PZT/EOM crossover yesterday. But I did measure the OLTF for various values of the FSS gains. At the current value of +20dBm, the PC drive signal is hovering around 1.5V. This bit of optimization needs to be done more systematically.
- I've edited mcup and mcdown to reflect the new gains.

Some general remarks
- The whole point of this exercise was to increase the modulation depth for the 29.5MHz signal.
- By my estimate, assuming 8mrad/V modulation index for the EOM and a gain of 0.6 at 29.5 MHz in the triple resonant box, we should have 100mrad of modulation after installing the amplifier (compared to 4mrad before the change).
- The actual RF power at 29.5 MHz at the input/output of the triple resonant box has not yet been measured.
- The WFS input error signal levels have to be re-measured (so I've turned off the inputs to the digital WFS filters for now)
|
|
|
|
|